Abner​ Corporation's bonds mature in 18 years and pay 8 percent interest annually. If you purchase the bonds for ​$1,050 what is your yield to​ maturity?
Answer #1

Yield To Maturity(YTM) = (interest per annum + average other cost per annum) / average fund employed

Average other cost per annum = (Redemption price - Current market price) / life remaining to maturity

= (1000-1050) / 18

= -50/18

= -2.77777777778

Average fund employed = (.4*Redemption price)+ (.6*Current market price)

= (.4*1000)+(.6*1050)

= 400+630

= 1030

Yield To Maturity(YTM) = ((1000*8%)-2.77777777778)/1030

= 77.2222222222/1030

= 7.49%